# Interpretable Vision-Language Survival Analysis with Ordinal Inductive Bias for Computational Pathology

### v3

@@ 6 Conclusion
This paper presents Vision-Language Survival Analysis (VLSA) for computational pathology. It is the first VL-based SA framework for histopathology WSIs. Different from current SA paradigm in CPATH, our VLSA proposes to leverage language-encoded prognostic priors as auxiliary signals to improve weakly-supervised WSI representation learning. Moreover, considering the intrinsic characteristics of SA, we propose ordinal survival prompt learning and ordinal incidence function prediction. Owing to the introducing of language-encoded prognostic priors, our VLSA enjoys a good interpretability, supported by our Shapley values-based interpretation method. The extensive experiments on five datasets verify the effectiveness of our scheme. On one hand, VLSA obtains new state-of-the-art performances in SA tasks; on the other hand, it often shows clear superiority over existing schemes in terms of few-shot learning and model efficiency. Notably, our comparative experiments and analyses suggest that vision-language-based learning paradigm can offer weakly-supervised MIL an effective means to learn valuable prognostic clues from gigapixel WSIs. This finding is likely to provide a new solution for improving the weak supervision in MIL, thereby paving a new way for survival analysis in computational pathology.
Appendix A Overview of Appendix
This appendix first provides related preliminary knowledge in Appendix B, including survival analysis (SA) (B.1), weakly-supervised MIL (B.2), Earth Mover’s Distance (EMD) (B.3), and Shapley value (B.4). In Appendix C, we further explain and discuss our two ordinal inductive biases to help readers better understand them. Then, we give more experimental details in Appendix D, including datasets (D.1), baselines (D.2), evaluation metrics (D.3), and the implementation details of VLSA (D.4). Finally, we present additional experimental results in Appendix E.
Appendix B Preliminaries
B.1 Survival Analysis
Survival analysis (SA), also known as time-to-event analysis, is one of the primary statistical approaches for analyzing data on time to event (Kaplan & Meier, 1958; Cox, 1975). Time-to-event data is usually denoted by $\mathcal{D}=\{{\bm{X}}_{i},t_{i},\delta_{i}\}_{i=1}^{N}$, where ${\bm{X}}_{i}$ represents individual characteristics and $\delta_{i}$ is the individual status of event of interest (e.g., death and machine failure) observed at time $t_{i}$. Particularly, $\delta=0$, called censorship in SA, indicate that event does not occur during follow-up observation. They are also considered in analysis and modeling. One common task in SA is to predict the risk of event occurrence. It covers a wide range of applications in healthcare and finance. Next, we give some key concepts in SA for reference.
@@ B.1 Survival Analysis
When prediction target is incidence function or hazard function, survival time is usually preprocessed as discrete time labels for analysis. We compare these two prediction targets and try different time discretization settings in our VLSA. The results can be found in Appendix E. Evaluation metrics for SA models are elaborated in Appendix D.3.
B.2 Weakly-Supervised Multi-Instance Learning
@@
B.3 Earth Mover’s Distance
@@ B.4 Shapley Value
Appendix C Further Discussion on the Ordinal Inductive Biases in VLSA
Our VLSA considers two ordinal inductive biases: i) ordinal survival prompts and ii) ordinal incidence function. For better understanding, next we first explain them in detail and then discuss the connection and difference between them.
@@ Appendix C Further Discussion on the Ordinal Inductive Biases in VLSA
as depicted in Figure 1(b). Our experiments (Table 3 and Table 8) confirm the effectiveness of these two ordinal inductive biases: i) ordinal survival prompts tend to enhance the model’s ability in risk discrimination; and ii) ordinal incidence function could often reduce the bias of predicted survival time (MAE).
@@
Figure 4: An example to illustrate the case that incidence function is not ordinal when prompts are. $L_{c,c+1}$ is the bisector of the angle between ${\bm{f}}_{\text{text}}^{c}$ and ${\bm{f}}_{\text{text}}^{c+1}$. When ${\bm{f}}_{\text{image}}$ falls into the gray area, $\text{cos}({\bm{f}}_{\text{image}},{\bm{f}}_{\text{text}}^{c+1})>\text{cos}({\bm{f}}_{\text{image}},{\bm{f}}_{\text{text}}^{c+2})$ does not hold.
Appendix D More Experimental Details
D.1 Datasets
@@
The preprocessing of histopathology WSIs contains the following key steps: i) tissue region segmentation, ii) image patching, and iii) patch feature extraction. Following CONCH (Lu et al., 2024), each WSI is divided into patches with size of 448 $\times$ 448 pixels at 20$\times$ magnification. In the last step, CONCH’s image encoder is used to obtain patch features. We utilize a toolkit provided by CLAM  [^1] Available at https://github.com/mahmoodlab/CLAM (Lu et al., 2021) for these steps. The statistical details on processed instances (patches) are provided in Table 6.
@@
Table 6: Dataset details. The last two rows are calculated at patient level. Maximum value in bold.
@@
D.2 Baselines
Vision-Only Baselines We compare the baselines originally proposed for WSI classification, i.e., ABMIL (Ilse et al., 2018), TransMIL (Shao et al., 2021b), and ILRA (Xiang & Zhang, 2023), and those for WSI survival analysis, i.e., DeepAttnMISL (Yao et al., 2020) and Patch-GCN (Chen et al., 2021). DeepAttnMISL and Patch-GCN are the most representative method based on cluster and graph, respectively.
Vision-Language Baselines Since the vast majority of existing works focus on classification or segmentation tasks, we adapt the related methods to WSI survival analysis. These methods include MI-Zero (Lu et al., 2023) for zero-shot WSI classification, CoOp (Zhou et al., 2022) for context prompt optimization, and OrdinalCLIP (Li et al., 2022) for VL-based ordinal regression.
@@ D.2 Baselines
  • ABMIL, TransMIL, and ILRA: we adapt them to SA by replacing their prediction target with incidence function and using the $\mathcal{L}_{\text{MLE}}$ same as our VLSA. Moreover, their time discretization setting (Section 3.2) is also the same as VLSA.
@@
  • MI-Zero, CoOp, and OrdinalCLIP: for MI-Zero, we try different score aggregation methods proposed by it and report one with the best performance, i.e., mean-based aggregation. Moreover, we tried different survival prompts for it and finally adopt the best: context = “the cancer prognosis reflected in the pathology image is”, classes = [“very poor”, “poor”, “good”, “very good”]. The implementation of CoOp and OrdinalCLIP are stated in the main paper. Unless specified, they use the same setting as our VLSA.
D.3 Evaluation Metrics
Our main evaluation metrics contain concordance index (CI), mean absolute error (MAE), and distribution calibration (D-cal). They are adopted to evaluate the performance of SA models in risk discrimination and distribution calibration. We utilize the toolkit from SurvivalEVAL  [^2] Available at https://github.com/shi-ang/SurvivalEVAL (Qi et al., 2023b) for performance evaluation. The section is adapted from Qi et al. (2023a). Readers could refer to Qi et al. (2023a) for more details. We show the details of these metrics below.
D.4 Implementation Details of VLSA
This section provides the implementation details of our VLSA, including the description of prognostic priors, survival prompt settings, hyper-parameter settings, and training settings. More details could be found in our source code available at https://github.com/liupei101/VLSA.
@@ D.4 Implementation Details of VLSA
Training Settings In model training, we set the number of epochs to 10, learning rate to 0.0002, batch size to 1 (one bag), the step of gradient accumulation to 32, and the optimizer to Adam with a weight decay rate of 0.00001. These setting are shared across five datasets, without dataset-specific fine-tuning. Please refer to our source code for more training details.
Appendix E Additional Experimental Results
@@
Effect of Ordinal Inductive Bias without Prognostic Priors in VLSA We show its results in Table 8. We can find that ordinal prompts obtain an improvement of around 1.3% in average CI when language-encoded prognostic priors are absent in VLSA. This improvement is larger than that with prognostic priors (Table 3). Moreover, we see the improvements in MAE when considering the ordinality of incidence function (Proba.) in optimization. This experiment could also confirm the effectiveness of our introduced ordinal inductive biases.
Table 8: Effect of ordinal inductive bias without prognostic priors in VLSA. ABMIL is used for multi-instance aggregation in this experiment.
Ordinality TCGA Average D-cal
Prompt Proba. BLCA BRCA GBMLGG LUAD UCEC CI MAE Count
0.5971 0.5994 0.7853 0.5750 0.6840 0.6482 28.70 5
($\pm$ 0.033) ($\pm$ 0.086) ($\pm$ 0.015) ($\pm$ 0.064) ($\pm$ 0.070)
0.6037 0.6202 0.7893 0.6053 0.6836 0.6604 28.01 5
($\pm$ 0.043) ($\pm$ 0.046) ($\pm$ 0.018) ($\pm$ 0.065) ($\pm$ 0.036)
0.5997 0.6049 0.7846 0.5818 0.6769 0.6496 26.99 3
($\pm$ 0.033) ($\pm$ 0.104) ($\pm$ 0.016) ($\pm$ 0.056) ($\pm$ 0.065)
0.6083 0.6180 0.7908 0.6048 0.6908
Different Time Discretization Settings Here we examine the impact of time discretization on model performance. Specifically, we try to cut survival time into 4 bins or using quantile time points. The results are shown in Table 9. We can see that quantile time points or 4 time bins could often lead to poor calibration in predictive survival distribution. A poorly-calibrated distribution often indicates over-confident survival predictions. Thus, we follow the setting of Haider et al. (2020) in time discretization, i.e., uniformly discretize time into $\sqrt{N_{e}}$ bins.
Table 9: Different time discretization settings. $N_{e}$ is the number of patients with event.
Time cutoff #Cuts TCGA Average D-cal
BLCA BRCA GBMLGG LUAD UCEC CI MAE Count
Quantile $\sqrt{N_{e}}$ 0.6255 0.6567 0.8006 0.6150 0.7300 0.6856 21.92 2
($\pm$ 0.015) ($\pm$ 0.064) ($\pm$ 0.013) ($\pm$ 0.031) ($\pm$ 0.038)
Uniform 4 0.6124 0.6458 0.7907 0.6312 0.7674 0.6895 23.25 1
($\pm$ 0.020) ($\pm$ 0.056) ($\pm$ 0.010) ($\pm$ 0.038) ($\pm$ 0.041)
0.6176 0.6652 0.8002 0.6370 0.7571
Comparison with Hazard-Based Survival Modeling Since individual hazard function is also utilized in some SA models, we further try to take it as the prediction target of VLSA. Following Zadeh & Schmid (2020), the settings of this experiment are as follows: i) survival time is discretized into 4 bins based on quantile time points; ii) we derive the prediction of individual hazard function by applying $\text{Sigmoid}(\cdot)$ to the scaled similarity scores, i.e., $\{\tau\cdot\text{cos}({\bm{f}}_{\text{image}},{\bm{f}}_{\text{text}}^{c})\}_{c=1}^{C}$; and iii) the MLE loss corresponding to hazard function prediction is adopted to optimize the model.
@@
PLIP as the Foundational VLM for Survival Analysis As another foundational VLM in computational pathology, PLIP (Huang et al., 2023) is pretrained on the pathology image-text pairs crawled from Twitter. It is the first work exploring VLMs for pathology, developed before CONCH (Lu et al., 2024). We further adopt it for VL-based methods in this experiment.

### v4

@@ 5 Conclusion
This paper presents a new Vision-Language Survival Analysis (VLSA) paradigm. It designs the first VL-based SA framework for histopathology WSIs. Different from current SA paradigm in CPATH, VLSA proposes to leverage language-encoded prognostic priors as auxiliary signals to improve weakly-supervised WSI representation learning. Moreover, considering the intrinsic characteristics of SA, VLSA proposes ordinal survival prompt learning and ordinal incidence function prediction. Owing to the introduce of language-encoded prognostic priors, our VLSA enjoys a good interpretability, supported by its Shapley values-based interpretation method. The extensive experiments on five datasets verify our scheme’s effectiveness: VLSA obtains new state-of-the-art performances in SA tasks and it could often show clear superiority over existing schemes in terms of data and computation efficiency. Notably, our empirical results suggest that VL-based learning paradigm can offer weakly-supervised MIL an effective means to learn valuable prognostic clues from gigapixel WSIs. This finding is likely to provide an alternative solution for improving the weak supervision in MIL, thereby paving a new way for survival analysis in computational pathology.
6 Acknowledgment
This work is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62476049, partly by the Fundamental Research Funds for the Central Universities under Grant No. ZYGX2022YGRH015 and the project of Sichuan Provincial Department of Science and Technology under Grant No. 2023YFQ0011. The authors would like to thank Mahmood Lab for making CONCH public to facilitate the research, and anonymous reviewers for their constructive comments to improve this work.
Appendix
[…]
Appendix A Preliminaries
A.1 Survival Analysis
Survival analysis (SA) Also known as time-to-event analysis, SA is one of the primary statistical approaches for analyzing data on time to event (Kaplan & Meier, 1958; Cox, 1975). Time-to-event data is usually denoted by $\mathcal{D}=\{{\bm{X}}_{i},t_{i},\delta_{i}\}_{i=1}^{N}$, where ${\bm{X}}_{i}$ represents individual characteristics and $\delta_{i}$ is the individual status of event of interest (e.g., death and machine failure) observed at time $t_{i}$. Particularly, $\delta=0$, called censorship in SA, indicate that event does not occur during follow-up observation. They are also considered in analysis and modeling. One common task in SA is to predict the risk of event occurrence. It covers a wide range of applications in healthcare and finance. Next, we give some key concepts in SA for reference.
@@ A.1 Survival Analysis
When prediction target is incidence function or hazard function, survival time is usually preprocessed as discrete time labels for analysis. We compare these two prediction targets and try different time discretization settings for our VLSA. The results can be found in Appendix D.6.
A.2 Weakly-Supervised MIL
@@
A.3 Earth Mover’s Distance
@@ A.4 Shapley Value
Appendix B More Discussions on Ordinal Inductive Bias
Our VLSA considers two ordinal inductive biases: i) ordinal survival prompts and ii) ordinal incidence function. For better understanding, in this section we first explain them in detail and then discuss the connection and difference between them.
@@ Appendix B More Discussions on Ordinal Inductive Bias
as depicted in Figure 1(b). Our experiments (Table 3) confirm the effectiveness of these two ordinal inductive biases: i) ordinal survival prompts tend to enhance the model’s ability in risk discrimination; and ii) ordinal incidence function could often reduce the bias of predicted survival time (MAE).
Figure 4: An example to illustrate the case that incidence function is not ordinal whereas prompts are. $L_{c,c+1}$ is the bisector of the angle between ${\bm{f}}_{\text{text}}^{c}$ and ${\bm{f}}_{\text{text}}^{c+1}$. When ${\bm{f}}_{\text{image}}$ falls into the gray area, $\text{cos}({\bm{f}}_{\text{image}},{\bm{f}}_{\text{text}}^{c+1})>\text{cos}({\bm{f}}_{\text{image}},{\bm{f}}_{\text{text}}^{c+2})$ does not hold.
@@
Appendix C More Experimental Details
C.1 Datasets
@@
The preprocessing of histopathology WSIs contains the following key steps: i) tissue region segmentation, ii) image patching, and iii) patch feature extraction. Following CONCH (Lu et al., 2024), each WSI is divided into patches with size of 448 $\times$ 448 pixels at 20$\times$ magnification. In the last step, CONCH’s image encoder is used to obtain patch features. We utilize a toolkit provided by CLAM  [^1] https://github.com/mahmoodlab/CLAM (Lu et al., 2021) for these steps. The statistical details on processed instances (patches) are provided in Table 5.
@@
Table 5: Dataset details. The last two rows are calculated at patient level.
@@
C.2 Baselines
Vision-Only Baselines We compare the baselines originally proposed for WSI classification, i.e., ABMIL (Ilse et al., 2018), TransMIL (Shao et al., 2021b), ILRA (Xiang & Zhang, 2023), and R2T-MIL (Tang et al., 2024), and those for WSI survival analysis, i.e., DeepAttnMISL (Yao et al., 2020) and Patch-GCN (Chen et al., 2021a). DeepAttnMISL and Patch-GCN are the most representative method based on cluster and graph, respectively.
[…]
@@ C.2 Baselines
  • ABMIL, TransMIL, ILRA, and R2T-MIL: we adapt them to SA by replacing their prediction target with incidence function and using the $\mathcal{L}_{\text{MLE}}$ same as our VLSA. Moreover, their time discretization settings (Section 3.2) ares the same as VLSA.
@@
  • MI-Zero: In vision-end, we try different score aggregation methods proposed by it to obtain WSI representations and report the best one, i.e., mean-based aggregation. In language-end, we try different survival prompts for MI-Zero and finally adopt the best: context = “the cancer prognosis reflected in the pathology image is”, classes = [“very poor”, “poor”, “good”, “very good”]. The final four survival prompts are encoded into four textual vectors by the CONCH’s text encoder. They are then used to derive the prediction of IF via Eq. (6), same as our VLSA.
[…]
  • CoOp and OrdinalCLIP: the implementation of CoOp and OrdinalCLIP are stated in the main paper. Unless specified, they use the same setting as our VLSA.
C.3 VLSA
This section provides the implementation details of our VLSA, including the description of prognostic priors, survival prompt settings, hyper-parameter settings, and training settings. More details could be found in our source code.
@@ C.3 VLSA
Training Settings In model training, we set the number of epochs to 10, learning rate to 0.0002, batch size to 1 (one bag), the step of gradient accumulation to 32, and the optimizer to Adam with a weight decay rate of 0.00001. These setting are shared across five datasets, without dataset-specific fine-tuning. Please refer to our source code  [^2] https://github.com/liupei101/VLSA for more training details. All experiments are run on a machine with two NVIDIA GeForce RTX 3090 GPUs (24G).
C.4 Evaluation Metrics
Our main evaluation metrics contain concordance index (CI), mean absolute error (MAE), and distribution calibration (D-cal). They are adopted to evaluate the performance of SA models in risk discrimination and distribution calibration. We utilize the toolkit from SurvivalEVAL  [^3] https://github.com/shi-ang/SurvivalEVAL (Qi et al., 2023b) for performance evaluation. The section is adapted from Qi et al. (2023a). Readers could refer to Qi et al. (2023a) for more details. We show the details of these metrics below.
C.5 Time Discretization Settings
This section provides the details of time discretization settings for all compared SA models. For discrete SA models, survival time is often required to be transformed into a discrete variable at first.
Conventional Settings As stated in Appendix A.1, discrete SA models are usually categorized into two groups: hazard-based and incidence-based. Their conventional settings in time discretization are as follows:
  • Hazard-based SA models: Most works (Chen et al., 2021b; Xu & Chen, 2023; Zhou & Chen, 2023; Zhang et al., 2024b) follow the settings used in Zadeh & Schmid (2020) and Chen et al. (2021a), i.e., using 4 time bins with quantiles as cutoff points.
  • Incidence-based SA models: As suggested in Haider et al. (2020), the number of time bins is determined by $\sqrt{N_{e}}$, which could often lead to better probability calibration in SA models. It is adopted in many works (Yu et al., 2011; Lee et al., 2018; Qi et al., 2023a; b). Particularly, DeepHit (Lee et al., 2018), as the first deep learning-based SA model with incidence function as prediction target, sets the number of time bins to $\sqrt{N_{e}}$ and the length of each time bin to be equal (i.e., time cutoff points are uniformly distributed).
Settings for Baselines and VLSA We follow conventional settings in time discretization:
  • For the hazard-based baseline model, Patch-GCN (Chen et al., 2021a), we follow its original setting to use 4 time bins with quantiles as cutoff points. We also examine its performance under the same number of time bins as VLSA for a fair comparison (Table 11).
  • For the baseline models not originally proposed for SA, we adopt the same discretization settings as our VLSA, i.e., $\sqrt{N_{e}}$ time bins with equal length.
  • For our incidence-based VLSA, we use $\sqrt{N_{e}}$ time bins with equal length, as stated in the main paper. Besides, we also evaluate the performance of VLSA using different time discretization settings. Please refer to Table 10 for the results.
Appendix D Additional Experimental Results
D.1 Risk Grouping and Kaplan-Meier Analysis
This experiment is conducted to examine whether our VLSA could discriminate the patients with high-risk and low-risk. Its results are shown in Figure 5. For each dataset, the median risk of the entire cohort is used as the cutoff for risk grouping. From the results in Figure 5, we can observe that VLSA can discriminate between high- and low-risk patients with a significant difference (P-Value $<0.05$ by a log-rank test).
Figure 5: Risk Grouping and Kaplan-Meier Analysis. All patients within each dataset are grouped into two risk groups: low-risk (blue) and high-risk (orange). Patients’ risk predictions are derived from VLSA. The median risk of the entire cohort is adopted as the cutoff for risk grouping.
D.2 Additional Visualization Results
@@
D.3 Analysis of Computation Efficiency
Table 7: Computation efficiency evaluation. A random WSI (from BLCA) with 20,265 instances is used to measure the # MACs in model inference.
Model # Params # MACs
V ABMIL 137.74K 2.68G
TransMIL 683.55K 14.23G
ILRA 3.16M 34.79G
R2T-MIL 2.44M 29.62G
DeepAttnMISL 329.22K 2.68G
Patch-GCN 363.91K 3.41G
[…]
CoOp 434.18K 2.68G
OrdinalCLIP 409.60K 2.68G
VLSA 284.16K 0.25G
In this experiment, we evaluate the computation efficiency of all models. The number of learnable parameters (# Params) and Multiply–Accumulate operations (# MACs) are taken as main evaluation metrics. From the results shown in Table 7, we can see that VLSA not only has fewer learnable parameters (284.16K) than all baseline models except for ABMIL but also is the most efficient one (0.25G) in inference cost. Our further analysis is as follows.
D.4 Comparison with Hazard-Based Survival Modeling
Since hazard function is also utilized in some SA models, we further try to take it as the prediction target of VLSA. Following Zadeh & Schmid (2020), the settings of hazard-based survival modeling for VLSA are as follows: i) survival time is discretized into 4 bins based on quantile time points; ii) we derive the prediction of individual hazard function by applying $\text{Sigmoid}(\cdot)$ to the scaled similarity scores, i.e., $\{\tau\cdot\text{cos}({\bm{f}}_{\text{image}},{\bm{f}}_{\text{text}}^{c})\}_{c=1}^{C}$; and iii) the MLE loss corresponding to hazard function prediction is adopted to optimize the model.
@@ D.4 Comparison with Hazard-Based Survival Modeling
D.5 Comparison with Multi-Modal Survival Analysis Methods
In this experiment, we compare our VLSA with multi-modal SA models. The textual description of prognostic priors are adopted in VLSA to provide extra auxiliary signals for weakly-supervised learning. An alternative approach to leveraging these textual priors is multi-modal representation learning, i.e., integrating and fusing visual and textual features. Accordingly, here we compare our proposed method with multi-modal SA models.
[…]
The results of this experiment are shown in Table 9. These results show that our VLSA could often perform better than the compared multi-modal SA models, with an improvement of 2.85% over the runner-up in average CI. This improvement could indicate the superiority of our scheme.
Table 9: Comparison with Multi-Modal Survival Analysis Models. These methods are originally proposed for learning multi-modal representations from WSIs and gene data. We replace their gene input with the language-encoded prognostic priors $E_{\text{text}}(\mathcal{T}_{\text{prog}})$ for comparisons.
Method TCGA Average D-cal
BLCA BRCA GBMLGG LUAD UCEC CI MAE Count
MCAT 0.5899 0.6322 0.7980 0.5768 0.6827 0.6559 31.44 3
($\pm$ 0.020) ($\pm$ 0.035) ($\pm$ 0.020) ($\pm$ 0.043) ($\pm$ 0.022)
[…]
($\pm$ 0.040) ($\pm$ 0.016) ($\pm$ 0.012) ($\pm$ 0.055) ($\pm$ 0.049)
[…]
($\pm$ 0.031) ($\pm$ 0.041) ($\pm$ 0.019) ($\pm$ 0.071) ($\pm$ 0.052)
MOTCat 0.6227 0.6301 0.7967 0.5617 0.7234 0.6669 29.73 3
($\pm$ 0.029) ($\pm$ 0.047) ($\pm$ 0.026) ($\pm$ 0.070) ($\pm$ 0.033)
0.6176 0.6652 0.8002 0.6370 0.7571
VLSA ($\pm$ 0.025) ($\pm$ 0.057) ($\pm$ 0.010) ($\pm$ 0.027) ($\pm$ 0.045) 0.6954 25.15 5
D.6 Different Time Discretization Settings
Different Time Discretization Settings for VLSA Here we examine the impact of time discretization on model performance. Specifically, we try to cut survival time into 4 bins or using quantile time points. The results are shown in Table 10. We can see that quantile time points or 4 time bins could often lead to poor calibration in predictive survival distribution. A poorly-calibrated distribution often indicates over-confident survival predictions. Thus, we follow the setting of Haider et al. (2020) in time discretization, i.e., uniformly discretize time into $\sqrt{N_{e}}$ bins.
Table 10: Different time discretization settings for VLSA. $N_{e}$ is the number of patients with event.
Time cutoff #Cuts TCGA Average D-cal
BLCA BRCA GBMLGG LUAD UCEC CI MAE Count
Quantile 4 0.6238 0.6483 0.7962 0.6240 0.7295 0.6844 24.40 0
($\pm$ 0.036) ($\pm$ 0.060) ($\pm$ 0.014) ($\pm$ 0.028) ($\pm$ 0.057)
Quantile $\sqrt{N_{e}}$ 0.6255 0.6567 0.8006 0.6150 0.7300 0.6856 21.92 2
($\pm$ 0.015) ($\pm$ 0.064) ($\pm$ 0.013) ($\pm$ 0.031) ($\pm$ 0.038)
Uniform 4 0.6124 0.6458 0.7907 0.6312 0.7674 0.6895 23.25 1
($\pm$ 0.020) ($\pm$ 0.056) ($\pm$ 0.010) ($\pm$ 0.038) ($\pm$ 0.041)
0.6176 0.6652 0.8002 0.6370 0.7571
Comparison with Patch-GCN under the Same Number of Time Bins We use the same number of time bins for Patch-GCN and VLSA to further compare their performance. The results are shown in Table 11. From these results, we can see that our VLSA could often outperform Patch-GCN in terms of average CI and MAE even when the number of time bins is set to the same value.
Table 11: Comparison with Patch-GCN under the same number of time bins.
#Cuts Method TCGA Average D-cal
BLCA BRCA GBMLGG LUAD UCEC CI MAE Count
Patch-GCN 0.6124 0.6375 0.7999 0.5922 0.7212 0.6726 26.70 2
($\pm$ 0.031) ($\pm$ 0.033) ($\pm$ 0.021) ($\pm$ 0.053) ($\pm$ 0.025)
0.6124 0.6458 0.7907 0.6312 0.7674
4 VLSA ($\pm$ 0.020) ($\pm$ 0.056) ($\pm$ 0.010) ($\pm$ 0.038) ($\pm$ 0.041) 0.6895 23.25 1
Patch-GCN 0.6054 0.6300 0.7890 0.5912 0.6967 0.6625 28.65 5
($\pm$ 0.049) ($\pm$ 0.068) ($\pm$ 0.020) ($\pm$ 0.029) ($\pm$ 0.030)
0.6176 0.6652 0.8002 0.6370 0.7571
$\sqrt{N_{e}}$ VLSA ($\pm$ 0.025) ($\pm$ 0.057) ($\pm$ 0.010) ($\pm$ 0.027) ($\pm$ 0.045) 0.6954 25.15 5
D.7 Adopting PLIP as VLM for Survival Analysis
As another foundational VLM in computational pathology, PLIP (Huang et al., 2023) is pretrained on the pathology image-text pairs crawled from Twitter. It is the first work exploring VLMs for pathology, developed before CONCH (Lu et al., 2024). We further adopt it for VL-based methods in this experiment.